OpenStudy (anonymous):

An object is thrown directly up (positive direction) with a velocity (vo) of 20.0 m/s and do= 0. Determine how long it takes to get to the maximum height of 24.0 m.

OpenStudy (anonymous):

at max height v= 0 so v^2 = u^2 - 2gs . v= 0 . so s = u^2/2g 20*20/2*10

OpenStudy (anonymous):

does it include a formula?

OpenStudy (anonymous):

yes its a formula where g = gravitational acceleration

OpenStudy (anonymous):

40/20 is 2 is that the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no 200 is the answer check the calculation

OpenStudy (anonymous):

`the answers i have that are close to 200 are 20.4 or 2.04

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

at highest point v=0 we know v=u - gt g=9.8 u=20 v=0 t=u/g=20/9.8=2.04

OpenStudy (anonymous):

u can also get by using formula s=u*t-.5*g*t^2 s=24 u=20 g=9.8 solve for t